# VGS-ATD: Robust Distributed Learning for Multi-Label Medical Image Classification Under Heterogeneous and Imbalanced Conditions

### v1

@@ V-D Robustness Analysis
  • Phase 1: All methods correctly predicted tuberculosis by focusing on the diseased region. For the second sample (COVID-19), FL and SL misclassified the image, focusing on irrelevant areas of bone and soft tissue. While CL correctly localized the lung region, VGS-ATD showed a more precise focus within the lung, avoiding attention to non-pathological structures.
  • Phase 2: With increased data heterogeneity, baseline methods began to show degradation. In the fracture sample, CL focused on background noise, and FL localized the bone but missed the fracture. Only SL and VGS-ATD correctly highlighted the fracture region. For the meningioma case, none of the baselines correctly localized the tumour. FL spread attention across the whole brain, while CL and SL failed to detect the tumor region. In contrast, VGS-ATD accurately identified the lesion.
  • Phase 3: VGS-ATD maintained reliable performance with new color medical images. In the melanoma and basophil cell samples, it accurately localized disease-specific features. Although CL and SL produced correct melanoma predictions, SL’s attention was unfocused, casting doubt on its reliability. Moreover, both baselines failed on the basophil cell sample, misdirecting attention to background regions.
@@
V-E Ablation Study
To assess the contribution of key components in VGS-ATD, we conducted a simple ablation study by sequentially removing four major modules: local transfer learning, node-to-node transfer learning, global feature selection, and soft voting. This study was conducted on a small group of nodes within the Phase 1 experiment setting to minimize computational costs. Table V presents the performance comparison of the global model when each component is excluded.
TABLE V: Ablation comparison on 3 nodes setting.
Settings Acc $\uparrow$ Pre $\uparrow$ Sensi $\uparrow$ F1 $\uparrow$ Without Base TL 81.42% 82.44% 74.89% 76.96% Without Node-Node TL 95.61% 94.72% 94.92% 94.77% Without Feature Selection 96.89% 95.78% 95.82% 95.79% Without Soft Voting 94.81% 93.86% 93.35% 93.48% VGS-ATD (Full) 97.33% 96.66% 97.24% 96.93% $\uparrow$: The higher the better, TL represent transfer learning.
The results show that removing any component results in reduced performance, confirming the effectiveness of each module. However, the degree of impact varies:
@@ V-E Ablation Study
  • Feature Selection has minimal effect in this setup, likely due to the limited complexity of feature combinations in the small 3-node setting. Its benefits might become more evident in large-scale environments with diverse feature spaces.
  • Soft Voting Ensemble improves robustness by aggregating predictions from multiple classifiers. Its absence reduces stability and overall accuracy, confirming the ensemble’s contribution.
@@ V-F Discussion
Based on extensive experiments and comparisons with three mainstream learning paradigms, VGS-ATD demonstrates superior overall performance, as shown in Table.VI. It consistently outperforms baselines by up to 17%, 22%, and 20% accuracy across the three experiment phases. From Phase 2 onward, it also achieves up to 90% reduction in training time and 50% savings in computational cost. Crucially, VGS-ATD mitigates catastrophic forgetting, with only a 1% drop in accuracy among the first three nodes after multi-phase expansion, whereas all baselines show substantial performance degradation on previously trained nodes. Its modular and scalable design allows seamless integration of new datasets and models without full retraining. In low-resource environments, its hierarchical structure enables efficient training by partitioning large datasets into manageable subsets for deployment on low-end devices. These strengths make VGS-ATD highly compatible with real-world clinical workflows and well-suited for continuous, non-disruptive learning.
@@
TABLE VI: Overall comparison of VGS-ATD and baselines.
@@
Phase Method Acc $\uparrow$ Time Cost $\downarrow$ Comp Cost $\downarrow$ CXR_Acc $\uparrow$ Phase1 CL 96.22 178.97 178.97 96.22 FL 83.15 76.64 185.98 83.15 SL 80.31 199.33 199.33 80.31 VGS-ATD 97.33 101.26 197.70 97.33 Phase2 CL 89.96 766.37 766.37 91.89 FL 72.35 86.64 637.57 74.35 SL 72.14 679.93 679.93 68.59 VGS-ATD 93.96 111.26 456.25 97.06 Phase3 CL 84.92 1441.96 1441.96 75.98 SL 72.99 1805.25 1805.25 73.61 VGS-ATD 92.64 154.18 969.89 96.17 $\uparrow$: The higher the better, $\downarrow$: The lower the better. Both Acc are calculated using percentage, where CXR represents 3 Phase1 nodes. Both computational and time costs are calculated using minutes.
@@
Despite its strengths, a few challenges remain. One is data scarcity at the node level. While we adopt the ViT model for its strong performance, ViT requires more training samples than CNNs. In some cases, CNNs outperform ViTs on small datasets due to better robustness. Although we applied data augmentation to help reduce overfitting, it can also increase computational cost and limit training speed. Another challenge lies in the lack of inter-node contrastive learning. The above experiment suggests that VGS-ATD performs well when datasets differ in modality or feature space. However, when several nodes each contain visually similar features but represent different classes (as seen in some skin disease datasets), distinguishing between classes becomes difficult. We assume this is likely due to the one-shot aggregation strategy applied in VGS-ATD, where nodes train independently without exposure to features from other labels. Merging their weights without global fine-tuning may cause semantic confusion.
VI Conclusion
@@
In this article, we proposed VGS-ATD, a Vision Transformer-based, generalized, and scalable distributed learning framework tailored for complex medical environments. The framework is designed to tackle key challenges in medical machine learning, including data privacy, data heterogeneity, class imbalance, computational efficiency, and scalability. We conducted comprehensive experiments to compare VGS-ATD with three widely used learning paradigms under increasingly complex clinical scenarios. Across all experiment phases, VGS-ATD demonstrated strong and consistent performance, maintaining over 92% accuracy even when handling an 80-class classification within a unified model, and significantly outperforming both centralized and decentralized baselines. These experimental results highlight VGS-ATD’s potential as a solution for next-generation medical AI, particularly in scenarios that require robustness, adaptability, and distributed collaboration.
@@
In the future, we plan to explore a wider variety of datasets and modalities to evaluate and enhance the generalization capability of VGS-ATD in broader clinical settings. Additionally, we aim to investigate alternative backbone architectures to provide users with greater flexibility in model design.

### v3

@@ 5.4 Robustness Analysis
  • Phase 1: All methods correctly predicted tuberculosis by focusing on the diseased region. For the second sample (COVID-19), FedLD and NonIID SL misclassified the image, focusing on irrelevant areas of bone and soft tissue. While CL correctly localized the lung region, VGS-ATD showed a more precise focus within the lung, avoiding attention to non-pathological structures.
  • Phase 2: At phase 2, baseline methods began to show misclassification. In the fracture sample, CL focused on background noise, and FedLD localized the bone but missed the fracture. Only NonIID SL and VGS-ATD correctly highlighted the fracture region. For the meningioma case, none of the baselines correctly localized the tumour. FedLD spread attention across the whole brain, while CL and NonIID SL failed to detect the tumor region. In contrast, VGS-ATD accurately identified the lesion.
  • Phase 3: VGS-ATD maintained reliable performance with new color medical images. In the melanoma and basophil cell samples, it accurately localized disease-specific features. Although CL and NonIID SL produced correct melanoma predictions, NonIID SL’s attention was unfocused, casting doubt on its reliability. Moreover, both baselines failed on the basophil cell sample, misdirecting attention to background regions.
@@
5.5 Ablation Study
To assess the contribution of key components in VGS-ATD, we conducted a simple ablation study by sequentially removing four major modules: local transfer learning, node-to-node transfer learning, PCA feature selection, and classifier ensemble using soft voting. This study was conducted on a small group of nodes within the Phase 1 experiment setting to minimize computational costs. Table 5 presents the performance comparison of the global model when each component is excluded.
Table 5: Ablation comparison on 3 nodes setting.
Settings Acc $\uparrow$ Pre $\uparrow$ Sensi $\uparrow$ F1 $\uparrow$ Without Base TL 81.42% 82.44% 74.89% 76.96% Without Node-Node TL 95.61% 94.72% 94.92% 94.77% Without PCA Selection 96.89% 95.78% 95.82% 95.79% Without Classifier Ensemble 94.81% 93.86% 93.35% 93.48% VGS-ATD (Full) 97.33% 96.66% 97.24% 96.93% $\uparrow$: The higher the better, TL represent transfer learning.
The results show that removing any component results in reduced performance, confirming the effectiveness of each module. However, the degree of impact varies as follow:
@@ 5.5 Ablation Study
  • PCA Feature Selection has minimal effect in this setup, likely due to the limited complexity of feature combinations in the small 3-node setting. Its benefits should become more evident in large-scale environments with diverse feature spaces.
  • Classifier Ensemble improves robustness by aggregating predictions from multiple classifiers. Its absence reduces stability and overall accuracy, confirming the ensemble’s contribution.
@@ 5.6 Discussion
Based on extensive experiments and comparisons with three mainstream learning paradigms, VGS-ATD demonstrates superior overall performance, as shown in Table.6, Its overall accuracy performance consistently outperforms state-of-the-art decentralized methods that are tailored for data heterogeneity by up to 17%, 21%, and 19% accuracy across the three experiment phases. Benefited from the one-time weight aggregation, the proposed VGS-ATD also achieves up to 90% reduction in training time and 50% savings in computational cost compared to baselines. Additionally, by comparing the accuracy performance change in the phase 1 chest x-ray dataset, the VGS-ATD has only 1% drop after multi-phase expansion, whereas all baselines show substantial performance degradation on previously trained nodes. The modular and scalable design of VGS-ATD allows seamless integration of new datasets and models without full retraining. In low-resource environments, its hierarchical structure enables efficient training by partitioning large datasets into manageable subsets for deployment on low-end devices. These strengths make VGS-ATD highly compatible with real-world clinical workflows and well-suited for continuous, non-disruptive learning.
@@
Table 6: Overall comparison of VGS-ATD against state-of-the-art baselines in Phase1 (3 Nodes with 4 classes), Phase2 (15 Nodes with 35 classes), and Phase3 (29 Nodes with 80 classes).
@@
Phase Method Acc $\uparrow$ CXR_Acc $\uparrow$ Time Cost $\downarrow$ Comp Cost $\downarrow$ Phase1 CL 96.22 96.22 178.97 178.97 FedLD [zeng2024tackling] 83.15 83.15 76.64 185.98 NonIID SL [gao2022new] 80.31 80.31 199.33 199.33 VGS-ATD (ours) 97.33 97.33 101.26 197.70 Phase2 CL 89.96 91.89 766.37 766.37 FedLD [zeng2024tackling] 72.74 75.84 86.64 637.57 NonIID SL [gao2022new] 72.14 77.00 679.93 679.93 VGS-ATD (ours) 93.96 97.06 111.26 456.25 Phase3 CL 84.92 75.98 1441.96 1441.96 NonIID SL [gao2022new] 73.45 75.04 1805.25 1805.25 VGS-ATD (ours) 92.64 96.17 154.18 969.89 $\uparrow$: The higher the better, $\downarrow$: The lower the better. Both Acc are calculated using percentage, where CXR_ACC represents average accuracy on three Phase1 chest x-ray nodes. Both computational and time costs are calculated using minutes.
@@
Despite its strengths, several challenges exist. One is data scarcity at the node level. While we adopt the ViT model for its strong performance, ViT requires more training samples than CNNs. In some cases, CNNs outperform ViTs on small datasets due to better robustness. Although we applied data augmentation to help reduce overfitting, it can also increase computational cost and limit training speed. Another issue is that we tested the whole system under an ideal environment without considering real-world communication overload. Due to the resource and time limitations, we plan to conduct verification experiments in the future. The last challenge lies in the lack of inter-node contrastive learning. The above experiment suggests that VGS-ATD performs well when datasets differ in modality or feature space. However, when several nodes each contain visually similar features but represent different classes (as seen in some skin disease datasets), distinguishing between classes becomes difficult. We assume this is likely due to the one-shot aggregation strategy applied in VGS-ATD, where nodes train independently without exposure to features from other labels. Merging their weights without global fine-tuning may cause semantic confusion.
6 Conclusion
@@
In this article, we propose VGS-ATD, a Vision Transformer based, generalized, and scalable distributed learning framework tailored for complex medical environments. The framework is designed to tackle key challenges in medical machine learning, including data privacy, data heterogeneity, class imbalance, computational efficiency, and scalability. We conducted comprehensive experiments to compare VGS-ATD with three widely used learning paradigms under increasingly complex clinical scenarios. Across all experiment phases, VGS-ATD demonstrated strong and consistent performance, maintaining over 92% accuracy even when handling an 80-class classification within a unified model, and significantly outperforming both centralized and decentralized baselines. These experimental results highlight VGS-ATD’s potential as a solution for next-generation medical AI, particularly in scenarios that require robustness, adaptability, and distributed collaboration.
@@
In the future, we plan to evaluate the stability of the whole system by testing the nodes with real-world communication overload scenarios, and exploring a wider variety of datasets and modalities to evaluate and enhance the generalization capability of VGS-ATD in broader clinical settings. Additionally, we aim to investigate alternative backbone architectures to provide users with greater flexibility in model design.
CRediT authorship contribution statement
Zehui Zhao:  Conceptualisation, Methodology, Validation, Writing – original draft, Laith Alzubaidi:  Conceptualisation, Methodology, Resources, Formal analysis, Writing – review & editing, Validation, Supervision, Project administration, Funding acquisition, Haider Alwzwazy:  Formal analysis, Writing – review & editing, Jinglan Zhang: Writing – review & editing, Formal analysis, Validation, Supervision, Clinton Fookes:  Writing – review & editing, Validation, Yuantong Gu:  Writing – review & editing, Validation, Funding acquisition.
Data statement
All the datasets used in this study are based on publicly available data, and the hyperlinks to these datasets are as follows: Tuberculosis Chest X-ray Dataset, Pneumonia Chest X-ray Dataset, Covid Chest X-ray Dataset, Fracture Spine Dataset, Fracture Body Dataset, Chest Covid CT Dataset, Kidney Cancer Dataset, Brain Tumor Dataset, PCOS Ultrasound Dataset, Liver CT Dataset, Brain Tumor CT Dataset, Liver Fibrosis Ultrasound Dataset, Breast Cancer MRI Dataset, Lumbar Coordinate Dataset, Breast Tumor MRI Dataset, Acute Lymphoblastic Leukemia Dataset, Breast Cancer Histopathological Dataset, Cervical Cancer Pathology Dataset, Lung and Colon Cancer Dataset, Lymphoma Cancer Dataset, Oral Cancer Dataset, Melanoma Cancer Dataset, Blood Cell Dataset, WCE Curated Colon Disease Dataset, Diabetic Foot Dataset, Retinal Cataract Dataset, Prostate Pathology Dataset, Otoscopic Dataset, and ORIGA Retinal Glaucoma Dataset.
Consent for publication
All authors have given their consent for publication.
Declaration of competing interest
All authors have declared that they have no conflict of interest.
Acknowledgements
I would like to thank my supervisors, colleagues, and reviewers for supporting my work and providing me with valuable feedback and suggestions.
Funding
Laith Alzubaidi and Yuantong Gu acknowledge the support received through the following funding schemes of the Australian Government: Australian Research Council (ARC) Industrial Transformation Training Centre (ITTC) for Joint Biomechanics under grant (IC190100020).